OpenStudy (anonymous):

prove : cos(x-pi/2)/sin(x-pi/2) = -tanx

OpenStudy (anonymous):

\[\frac{ \cos(x-\frac{ \pi }{ 2} )}{ \sin(x-\frac{ \pi }{ 2 }) } = -tanx\]

OpenStudy (anonymous):

I used cofunction identities to get sinx/cosx but that equals positive tan not negative tan, thats where im lost

OpenStudy (mathmale):

I'd strongly suggest that you use the difference formulas for the sine and cosine to evaluate \[\frac{ \cos(x-\frac{ \pi }{ 2} )}{ \sin(x-\frac{ \pi }{ 2 }) } \]

OpenStudy (mathmale):

For example: cos (a-b) = cos a cos b + sin a sin b

OpenStudy (mathstudent55):